OpenStudy (anonymous):

Solve each expression using the imaginary unit I. 71.sqrt-2-3 Solve each equation. 81.5x^2-2x+1=0 Solve each system by substitution 85.y=-x^2-7x-6 y=x^2-8x-12

OpenStudy (anonymous):

@iGreen can you just help me with number 71?

OpenStudy (igreen):

?? Is it this? \(\sqrt{-2} - 3\) or \(\sqrt{-2-3}\)

OpenStudy (anonymous):

The first one.

OpenStudy (igreen):

Okay.

OpenStudy (igreen):

Wait, so do they like want it in decimal form or..? Do you have choices?

OpenStudy (anonymous):

Ugh, it wants me to simplify each expression. . No I don't, it's a study guide.

OpenStudy (anonymous):

iGreen?

OpenStudy (mathmath333):

.\(\large\tt \begin{align} \color{black}{\sqrt{-2}-3\\~\\ }\end{align}\) is this is it?

OpenStudy (anonymous):

Yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\sqrt{-2}-3\\~\\ =\sqrt{2\times -1}-3\\~\\ =\sqrt{2}\times \sqrt{-1}-3\\~\\ =\sqrt{2}\times i-3-----------\color{blue}{(i=\sqrt{-1})}\\~\\ =\sqrt{2}i-3\\~\\}\end{align}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5x^2-2x+1=0\\~\\ \\~\\}\end{align}\) compare it with \(\large\tt \begin{align} \color{black}{ax^2+bx+c=0\\ }\end{align}\) so here \(\large\tt \begin{align} \color{black}{a=5,b=-2,c=1}\end{align}\) now apply formula \(\large\tt \begin{align} \color{black}{x=\dfrac{-b \pm \sqrt{b^2-4ac}}{2a}\\~\\ x=\dfrac{-(-2)\pm \sqrt{(-2)^2-4\times 5\times 1}}{2\times 5}\\~\\ =\dfrac{2\pm \sqrt{-16}}{10}\\~\\ =\dfrac{2\pm \sqrt{16}\times \sqrt{-1}}{10}\\~\\ =\dfrac{2\pm 4\times i}{10}\\~\\ =\dfrac{1+ 2i}{5}~~or~~\dfrac{1- 2i}{5}}\end{align}\)

OpenStudy (anonymous):

Thank you. xD

OpenStudy (mathmath333):

try the third